The sides of a rectangle are in the ratio of 3:2. What is the length of each side if the perimeter of the rectangle is 55 cm?
Answer by oberobic(2304)

Perimeter = 2*Length + 2*Width = 2L+2W = 2(L+W)
P = 55 cm
.
The tricky part here is to model the ratio of L:W = 3:2
.
We could try:
L/W = 3/2
.
L = 3/2*W
.
Substitute for L in the perimeter equation.
.
2(3/2*W + W) = 55
.
Multiply through
.
3W + 2W = 55
.
Collect
5W = 55
.
Divide both sides by 5
.
W = 11
.
L = 3/2*W
L = 33/2
L = 16.5
.
Check perimeter.
2*16.5 = 33
2*11 = 22
Total = 55
.
Is the ratio 3:2?
Yes.
16.5/11 = 1.5 = 3/2
.
Re-read question to see what the answer should be:
What is the length?
.
The length = 16.5 cm.
.
Done.